OpenStudy (anonymous):

What is a equation for a circle with center if (-2,3) and radius of 5?

OpenStudy (superdavesuper):

General eqn for a circle center at (A,B) and radius R is (x-A)^2 + (y-B)^2 = R^2

OpenStudy (anonymous):

what is a & b for?

OpenStudy (superdavesuper):

a n b are the x n y coordinates of the center

OpenStudy (anonymous):

okay how do I plus in the numbers bc I have radius?

OpenStudy (superdavesuper):

write down what you have and we will go from there :)

OpenStudy (anonymous):

(x+2)^2 + (y-3)^2= 5^2

OpenStudy (superdavesuper):

yup yup good job, katie! :) now 5^2=5x5=?

OpenStudy (anonymous):

25?

OpenStudy (superdavesuper):

yup yup :)